$SPAD/src/input exdiff.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool exdiff.output
)set message test on
)set message auto off
)clear all
--S 1 of 10
differentiate(sin(x) * exp(x**2),x)
--R
--R
--R              2                  2
--R             x                  x
--R   (1)  2x %e  sin(x) + cos(x)%e
--R                                                    Type: Expression(Integer)
--E 1
-- Input for page ExDiffSeveralVariables
)clear all
--S 2 of 10
differentiate(sin(x) * tan(y)/(x**2 + y**2),x)
--R
--R
--R                         2    2
--R        (- 2x sin(x) + (y  + x )cos(x))tan(y)
--R   (1)  -------------------------------------
--R                    4     2 2    4
--R                   y  + 2x y  + x
--R                                                    Type: Expression(Integer)
--E 2
--S 3 of 10
differentiate(sin(x) * tan(y)/(x**2 + y**2),y)
--R
--R
--R          2    2             2                       2    2
--R        (y  + x )sin(x)tan(y)  - 2y sin(x)tan(y) + (y  + x )sin(x)
--R   (2)  ----------------------------------------------------------
--R                               4     2 2    4
--R                              y  + 2x y  + x

--R                                                    Type: Expression(Integer)
--E 3
-- Input for page ExDiffMultipleI
)clear all
--S 4 of 10
differentiate(sin(x)/(x**2 + y**2),[x,y])
--R
--R
--R                           3     2
--R        8x y sin(x) + (- 2y  - 2x y)cos(x)
--R   (1)  ----------------------------------
--R               6     2 4     4 2    6
--R              y  + 3x y  + 3x y  + x
--R                                                    Type: Expression(Integer)
--E 4
--S 5 of 10
differentiate(sin(x)/(x**2 + y**2),[x,y,y])
--R
--R
--R                2     3             4     2 2     4
--R        (- 40x y  + 8x )sin(x) + (6y  + 4x y  - 2x )cos(x)
--R   (2)  --------------------------------------------------
--R                   8     2 6     4 4     6 2    8
--R                  y  + 4x y  + 6x y  + 4x y  + x
--R                                                    Type: Expression(Integer)
--E 5

-- Input for page ExDiffMultipleII
)clear all
--S 6 of 10
differentiate(cos(z)/(x**2 + y**3),[x,y,z],[1,2,3])
--R
--R
--R                    4      3
--R            (- 84x y  + 24x y)sin(z)
--R   (1)  --------------------------------
--R         12     2 9     4 6     6 3    8
--R        y   + 4x y  + 6x y  + 4x y  + x
--R                                                    Type: Expression(Integer)
--E 6
-- Input for page ExDiffHigherOrder
)clear all
--S 7 of 10

differentiate(exp(x**2),x,4)
--R
--R
--R                             2
--R            4      2        x
--R   (1)  (16x  + 48x  + 12)%e
--R                                                    Type: Expression(Integer)
--E 7
-- Input for page ExDiffFormalIntegral
)clear all
--S 8 of 10
f := integrate(sqrt(1 + t**3),t)
--R
--R
--R           t  +-------+
--R         ++   |  3
--R   (1)   |   \|%M  + 1 d%M
--R        ++
--R                                         Type: Union(Expression(Integer),...)
--E 8
--S 9 of 10
differentiate(f,t)
--R
--R
--R         +------+
--R         | 3
--R   (2)  \|t  + 1
--R                                                    Type: Expression(Integer)
--E 9
--S 10 of 10
differentiate(f * t**2,t)
--R
--R
--R             t  +-------+          +------+
--R           ++   |  3             2 | 3
--R   (3)  2t |   \|%M  + 1 d%M  + t \|t  + 1
--R          ++
--R                                                    Type: Expression(Integer)
--E 10
)spool
)lisp (bye)

———-

References^{[1] nothing}

6